Angles of a triangle. The second angle of a triangular
building lot is three times as large as the first. The
third angle is 10 degrees more than the sum of the other two
angles. Find the measure of the third angle.

Answer by rfer(16322)

x+3x+(x+3x+10)=180
8x+10=180
8x=170
x=21.25 degrees
3x=63.75 degrees
x+3x+10=95 degrees